Advanced Data Analysis, II

36-402  Homework 2  Spring, 1996

Homework due — In class, Thursday February 22.

The following problems are taken from the book “Applied Linear Statistical Models” by Neter, Wasserman, and Kutner. The data are available in the directory /afs/andrew/hss/stat/data/402.

1. Eye contact effect. In a study of the effect of eye contact (factor A) and gender of personnel officer(factor B) on a personnel officer’s assessment of likely job success of an application, 10 male and 10 female personnel officers were shown a frontview photograph of an applicant’s face and were asked to give the person in the photograph a success rating on a scale of 0 (tot failure) to 20 (outstanding success). Half of the officers in each gender group were chosen at random to receive a version of the photograph in which the subject made eye contact with the camera lens. The other half received a version in which there was no eye contact. The data are in eyecontact.dat The success ratings follow.

Gender of officer

Eye contact Male  Female
Present 11 15
7  12
12  14
6  11
10  16
Absent  12  14
16  17
10  13
13  20
14  18

(a) Analyze the data, looking at fitted values, residuals, the ANOVA table, and assumptions.

(b) Who would you rather have interview you? Is Eye contact important?

2. Hay fever relief. A research laboratory was developing a new compound for the relief of severe case of hat fever. In an experiment with 36 volunteers, the amounts of the two active ingredients in the compound were varied at three levels each. Randomization was used in assigning four volunteers to each of the nine treatments. The data on hours of relief follow, and are in hayfever.dat

Ingredient 2

Ingredient 1 Low Medium High
Low  2.4  4.6  4.8
2.7  4.2  4.5
2.3  4.9  4.4
2.5  4.7  4.6
Medium  5.8  8.9  9.1
5.2  9.1  9.3
5.5  8.7  8.7
5.3  9.0  9.4
High  6.1  9.9 13.5
5.7  10.5 13.0
5.9  10.6 13.3
6.2  10.1 13.2

(a) Analyze the data, looking at fitted values, residuals, the ANOVA table, and assumptions.

(b) Who would you rather have interview you? Is Eye contact important?

(c) Summarize the relationship between the factor levels and the response.

(d) What factor level(s) is/are best?

3. Disk Drive Service. The staff of a service center for electronic equipment includes three technicians who specialize in repairing three widely used makes of disk drives for desk-top computers. It was desired to study the effects of technician and make of disk drive on the service time. The data that follow show the number of minutes required to complete the repair job in a study where each technician was randomly assigned to five jobs on each make of disk drive. The data are in diskdrive.dat

Make of Drive

Technician Make 1 Make 2 Make 3
Technician 1  62  57  59
48  45  53
63  39  67
57  54  66
69  44  47
Technician 2  51  61  55
57  58  58
45  70  50
50  66  69
39  51  49
Technician 3  59  58  47
65  63  56
55  70  51
52  53  44
70  60  50

(a) Analyze the data, looking at fitted values, residuals, the ANOVA table, and assumptions.

(b) Who would you rather have repair your hardware? Does it matter what type of drive you have? Is there an interaction?

Let me know how things are going!